JB HiFi (JBH)

Instalment per

month ($’000s)

$32.9

Annual total

revenue ($millions)1

$7,095.3

Annual growth

in total revenue2

16.14%

Loan A (APR,

compounding frequency)

4.62%, semi- annually

Loan B (APR,

compounding

frequency)

4.52%,

quarterly

Loan C (APR,

compounding frequency)

4.24%, daily3

Property cost

$691,000

7 year bond

annual coupon rate

5.85%

7 year bond

current price

$99.95

5 year bond required rate of

return

4.85%

TVM and bond valuation questions (1 mark each):

a.            Your case company has just made a large sale on an instalment contract. The contract requires the customer to pay your case company the amount shown in Table 1 (JBH)at the end of every month for 5 years. Your case company would like the cash now for an investment and so has asked its bank to discount the instalment contract and pay it the discounted value. The bank will discount the contract at 6% APR, compounded monthly. How much cash will your case company receive from the bank?

b.            Your company has annual operating revenue as shown in Table 1. Assume this revenue will grow continuously at the annual rate shown in Table 1. What is your prediction for annual operating revenue in 7 years?

c.            Your company needs to borrow funds and has several options available to it, Loans A, B and

C. The interest rates (APR) for these options are given in Table 1. What is the EAR of the loan option the company should choose?

d.            Your company is buying new property for the amount given in Table 1. To finance this, the company’s bank has offered an amortised loan at 3.65% APR, quarterly compounding, with 11 years of quarterly payments. What quarterly payment will the company have to make on this loan? Assume that the entire property cost is financed and that payments are made at the end of each period.

e.            Your company has an issue of $100 par value annual coupon bonds with 7 years remaining until maturity. The annual coupon rate is given in Table 1, along with the current price of the bonds. What is the yield to maturity on the bonds?

f.            Your company has an issue of $1,000 par value bonds that offer a 8% coupon rate paid semi- annually. The bonds have 5 years remaining until maturity. The market’s required return on these bonds is given in Table 1. What is the amount of each coupon payment?
Answer #1

Part (a): The amount received on discounting installments is the present value of annuity arrived at $17,017,709.49 as follows:

1 Present Value of Annuity Payments at the end of each period 2 3 Present value of annuity is calculated using the formula PV

Part (b):Annual operating income in 7 years is the future value of current operating income, arrived at $ 20,222.90 Million as follows:

1 Future Value 2 3 Future value of an amount is calculated using the formula FV=P(1+r)^n 4 Where P= Principal (Presen Value),

Part (c):

EAR of various APRs given are as follows:

APR and compounding frequency

EAR
4.62% Semi annually 4.673361%
4.52% Quarterly 4.597193%
4.24% Daily 4.330915%

Details as follows:

R F4 fx =1*(1+F1/F3)^F3 B C 1 APR 2 Frequency of compounding 3 Number of times compounded a yeart 4 Future value of $ 1 FV =1

Part (d):

Quarterly payment on the loan of $691,000 (cost of property is $ 19,138.23 as follows:

8 A B C D 1 Equated Quarterly Installments (EQI) Payments at the end of each period 3 EQI is calculated using the formula EQI

Part (e); YTM is calculated at 5.87675266% using the RATE function of Excel as follows:

D Function FV $100 5.85% Annual Assumed for =RATE(E7,E8,29*-1,E2) C 1 Formula Cell ref. 2 Face Value (FV) Given 3 Coupon rate

Part (f) Amount of coupon payment is Par Value* Coupon Rate/ No. of payments a year.

In the given case,

Par Value= $1,000. Coupon Rate= 8%. Coupon payment frequency= Semi annual. Hence number of payments a year=2.

Therefore, amount of each coupon payment= $1,000*8%/2 = $1,000*4% = $40